# Decentralization of Ethereum’s Builder Market

### v1

@@ 5.1. Inequality in Block-building Capacity
In the context of auctions generally, the true value of the item being auctioned refers to its actual worth to the bidder. Recall that in MEV-Boost auctions (Sec. 2), a builder submits a block $B$ (a list of transactions) and a bid value $BV$ that the builder is willing to pay to the proposer if $B$ is proposed. The true value of $B$, denoted $TV(B)$, is typically higher than $BV$. One can think of $TV$ as the builder’s revenue and $BV$ as the cost. The difference $EV:=TV-BV$ is the profit pocketed by the builder, which we call extracted value (EV) — i.e., $EV(B)$ is the difference of the builder’s ETH balance after executing transactions in $B$. For example, suppose executing $B$ yields a net profit of $0.1$ ETH after paying $0.9$ ETH in bid value in the auction, then the true value (the revenue) is $1.0$ ETH.
@@
To compute $TV(B)$ of a given block $B$, we follow the standard approach in open source MEV analysis tools (EigenPhi, 2023; Flashbots, 2023b) and previous works (Heimbach et al., 2023; Thiery, 2023), though we built our own analysis toolchain and optimized its performance to handle the large volume of auction data. We compute $EV(B)$ using a modified Rust Ethereum execution client (Reth (paradigmxyz, 2024)) to execute transactions in $B$ and compute the net ETH balance changes of the bidder’s addresses. Most builders’ addresses are well-known partly because they need to build up a reputation. We cross-validated it with other sources to ensure that our dataset was the most complete one to date (see Sec. 3.1 for details.) Then, with bid value $BV(B)$ from the auction dataset, the true value $TV(B)$ is the sum of $EV(B)$ and $BV(B)$.
@@
To quantify the disparity of block-building capacity, we compute the quartile coefficient of dispersion (QCD) (Wikipedia, 2022) of the highest true values in a given slot. Specifically, suppose the true values of different builders in slot $s$ are $\vec{tv}_{s}$. Let $Q_{1}(\vec{x})$ and $Q_{3}(\vec{x})$ representing the first and third quartiles of $\vec{x}$ (i.e., 25% and 75% percentiles). The QCD of slot $s$ is
QCD is a real value in $[0,1]$. A larger QCD indicates greater dispersion (the intuition is that in a dataset with big dispersion, $Q_{3}$ is much larger than $Q_{1}$, thus $QCD$ approaches $1$). Since it uses quartiles, QCD has the benefit of being robust to outliers (Wikipedia, 2022).
Evaluation results.   We categorize the builders into three groups based on their market share from April to August 2023: top builders (top 5), middle builders (6-15), and tail builders (16-25). To understand how the magnitude of MEV in a slot affects the block-building capability, we categorize every slot into one of three tiers according to the final winning bid (as a proxy to the significance of the MEV opportunity in that slot): low MEV (below 0.0225 ETH, the 10th percentile of the winning bid values), high MEV (above 0.4586 ETH, 95th percentile) and medium MEV in between.
In Fig. 7, we plot the distribution of the QCD for MEV-Boost auctions from April 2023 to August 2023 for different builder groups and MEV tiers. (We sample the first week of each month. See Sec. 3.1 for details.) It reveals two trends. In each MEV tier, there is a clear increase in inequality (a high QCD means high inequality) from the top to the middle and then to the tail. That is, top builders have comparable block-building capabilities, but middle and tail builders have significantly different capabilities. This might be good news: top builders having similar abilities means they can meaningfully compete, which is positive for the decentralization of builder market.
Second, the inequality worsens as the MEV tier increases. One possible reason is that private order flow providers may tend to give significant MEV opportunities to select builders due to the trust barrier.
Figure 7. Violin graphs for the distribution of the QCD across three MEV tiers, for top, middle, and tail builders. The QCD measures the inequality of block-build capacity across builders. This figure reveals that such inequality worsens as MEV increases. Within a given MEV tier, inequality is greatest among tail builders, followed by middle builders, and is least among top builders.
Finding V: The inequality in block building is lowest among top builders, higher among middle builders, and most pronounced among tail builders, and the inequality worsens as a slot’s MEV increases.
5.2. Quantifying Competitiveness and Efficiency
The previous section shows that top-5 builders have comparable true values in most cases. Ideally, revenue from MEV-Boost auctions should be near optimal due to competition. However, whether real-world auctions induce such near-perfect competition is complicated by several confounding factors: builders may not be able to respond to bids due to network latency, they may collude implicitly or explicitly, or they may have incentives to overbid (e.g., to gain market share). In this section, we quantify the competitiveness and efficiency of the MEV-Boost auctions using what we call competitive index (CI) and efficient index (EI).
@@ 5.2. Quantifying Competitiveness and Efficiency
Namely, $CI(s)$ measures the relative difference between the winning bid and the second-highest true value. $CI(s)\geq 0$ indicates that the winning bid value is not less than the second-highest true value, satisfying the definition of a competitive auction. Conversely, $CI(s)<0$ means that the competition is insufficient.
@@
Figure 8. The distribution of CI and EI, where each dot represents an auction. The color represents density when multiple dots overlap. 88.84% of the MEV-Boost auctions are competitive (sum of the slots having $CI\geq 0$), and 79.74% of them are efficient (sum of the slots with $EI>0$). Dots on the diagonal line ($CI=EI$) are auctions where the winners bid the true values. The dots above this line are auctions where the winners bid lower than true values.
@@
Fig. 8 shows the distribution of the calculated CI and EI across all MEV-Boost auctions in the ultra sound relay dataset. Notably, 74.83% of the auctions are both competitive and efficient: the winner possesses the highest true value and bids higher than or equal to the second-highest true value. 88.84% (sum of the areas with $CI\geq 0$) of the auctions are competitive, yet only 79.74% (sum of the areas with $EI>0$) are efficient.
5.2.1. Implications of uncompetitive auctions
@@
We first note Fig. 8 shows a somewhat strict definition of competitiveness (i.e., $CI\geq 0$). In reality, builders may bid slightly lower than their true value to ensure profits. Therefore, it is reasonable to relax and consider an auction with $CI\geq-\delta$ to be competitive for a small $\delta$. We call this notion $\delta$-competitive. As shown in Fig. 14, when we consider $\delta=1\%$ or $\delta=2\%$, the percentage of uncompetitive auctions reduces to about 6% and 5%, respectively. That is, under the slightly relaxed definition, approximately 95% auctions were competitive.
Quantifying proposers’ losses.   The concern with uncompetitive auctions lies in the potential loss to validators. If such losses are significant, top validators might be motivated to operate their own builders. Furthermore, vertical integration between top proposers and builders could adversely impact the decentralization of the validators (Bahrani et al., 2024). Thus, quantifying the exact loss validators incur due to insufficient competition is crucial.
For each uncompetitive auction, we consider the second-highest true value as the ideal winning bid value and calculate the difference between this ideal value and the actual winning bid value. This difference represents the loss a validator incurs, which they would have received had the auction been competitive.
Table 3. Validators’ losses due to insufficient competition
Time #Slots #Slots w/ CI ¡ 0 (%) Profits (ETH) Losses (ETH) (%)
April 9-15 28,376 1,437 (5.06%) 2,711.32 47.17 (2.02%)
May 1-7 30,279 8,449 (27.90%) 9,367.50 118.50 (1.36%)
June 1-7 35,414 3,137 (8.86%) 4,346.51 24.72 (0.62%)
July 1-7 36,032 2,051 (5.69%) 3,945.67 18.41 (0.48%)
August 1-7 17,825 1,424 (7.99%) 2,171.07 12.29 (0.58%)
As shown in Tab. 3, we note that the total losses caused by uncompetitive auctions amount to 221.09 ETH in our study, which represents 0.98% of the total gains of proposers, with an average loss of 0.0134 ETH per uncompetitive auction.
The losses are particularly significant during May 1-7, 2023, when a total of 118.50 ETH was lost due to insufficient competition, which is 1.27% of their total gain. These losses may closely correlate with MEV opportunities, given that proposers’ total profits were also notably high during this period. Consequently, even though the losses seem to decrease, this reduction should not be interpreted as an indicator of increasingly intense competition. The concern regarding potential centralization persists, given that the losses experienced by validators have not been fully mitigated.
Finding VI: About 11% of the MEV-Boost auctions in our study are uncompetitive, resulting in total losses of 221.09 ETH for the proposers. These losses represent 0.98% of the total gains of proposers.
5.2.2. Reasons of inefficiency
20.26% of MEV-Boost auctions were not efficient. Further investigation is required to understand why. In an inefficient auction, the builder with the highest true value does not win the auction. There are two possible reasons. First, block subsidization allows a builder to win the auction without possessing the highest true value. Second, builders may “shade” their bids (bid shading means strategically placing bids slightly below one’s true value); if the builder with the highest true value “shades too much,” it will lose.
We analyzed the underlying reasons for all inefficient auctions. In 51.5% of cases, winners win the auctions due to subsidization. In 52.6% of cases, the builder with the highest true value shaded too much. It is important to note that auctions may exhibit inefficiency for multiple reasons. For instance, the winner may subsidize her bid while simultaneously, the builder with the highest true value shades. Thus, the cumulative percentage exceeds 100%.
It is important to note that inefficiency does not necessarily disadvantage the validators. In fact, in inefficient scenarios involving subsidization, validators may receive more value.
Finding VII: MEV-Boost auctions are generally competitive, yet their efficiency falls short of expectations. Further analysis reveals that 51.5% of inefficiencies are caused by subsidization, which, interestingly, does not disadvantage proposers but leads to increased earnings for them in such scenarios.
5.2.3. MEV’s implication on auction outcomes
@@
To further understand how the magnitude of MEV in a slot affects the competitiveness and efficiency of the auction, we follow the same methodology and categorize the auctions into three tiers: low MEV, medium MEV, and high MEV. The ratio of competitive and efficient auctions in each tier is shown in Fig. 9.
We note that competitiveness decreases as the MEV of a slot increases, which is consistent with the finding in Sec. 5.1 about the inequality of block-building capabilities in high-MEV slots.
The trend of efficiency is less expected. First, only 71.7% of the auctions within the low MEV tier are efficient. This can be attributed to the generally low overall bid values, rendering the auction outcomes vulnerable to alterations from minor subsidies. Second, while auctions tend to be more efficient in the medium MEV tier, we also observe that efficiencies worsen in the high MEV tier: only 77.2% of these auctions are efficient, and about 90% of inefficient cases result from bid shading, where the builders with the highest true values place their bids much lower than their true values.
@@
Figure 9. The ratio of auctions that are competitive and efficient across various MEV tiers.
Finding VIII: The competitiveness of MEV-Boost auctions deteriorates as a slot’s MEV increases. Efficiency may decline due to block subsidization in the low MEV tier and bid shading in the high MEV tier.
6. Related Work

### v3

@@ 4.1 Competitiveness of MEV-Boost Auctions
True value.   In the context of auctions generally, the true value of the item being auctioned refers to its actual worth to the bidder. Recall that in MEV-Boost auctions (Sec. 2), a builder submits a block $B$ (a list of transactions) and a bid value $BV$ that the builder is willing to pay to the proposer if $B$ is proposed. The true value of $B$, denoted $TV(B)$, is typically higher than $BV$. One can think of $TV$ as the builder’s revenue and $BV$ as the cost. The difference $EV:=TV-BV$ is the profit pocketed by the builder, which we call extracted value (EV) — i.e., $EV(B)$ is the difference of the builder’s ETH balance after executing transactions in $B$. For example, suppose executing $B$ yields a net profit of $0.1$ ETH after paying $0.9$ ETH in bid value in the auction, then the true value (the revenue) is $1.0$ ETH.
@@
To compute $TV(B)$ of a given block $B$, we follow the standard approach in open source MEV analysis tools [20, 31] and previous works [41, 80], though we built our own analysis toolchain and optimized its performance to handle the large volume of auction data. We compute $EV(B)$ using a modified Ethereum execution client (Reth [68]) to execute transactions in $B$ and compute the net ETH balance changes of the bidder’s addresses. Most builders’ addresses are well-known partly because they need to build up a reputation. We cross-validated it with other sources to ensure that our dataset was the most complete one to date (see Sec. 3.1 for details.) Then, with bid value $BV(B)$ from the auction dataset, the true value $TV(B)$ is the sum of $EV(B)$ and $BV(B)$.
@@
Namely, $CI(s)$ measures the relative difference between the winning bid and the second-highest true value. $CI(s)\geq 0$ indicates that the winning bid value is not less than the second-highest true value, satisfying the definition of a competitive auction, while $CI(s)<0$ means insufficient competition.
@@
Figure 4: The distribution of CI and EI, where each dot represents an auction. The color represents density when multiple dots overlap. 91.8% of the MEV-Boost auctions are competitive (sum of the slots having $CI\geq 0$), and 75.1% of them are efficient (sum of the slots with $EI>0$). Dots on the diagonal line ($CI=EI$) are auctions where the winners bid the true values. The dots above this line are auctions where the winners bid lower than true values.
@@
Analysis of uncompetitiveness and inefficiency.   Fig. 4 shows the calculated CI and EI distribution across all MEV-Boost auctions in the ultra sound relay dataset. Notably, 91.8% (sum of the areas with $CI\geq 0$) of the auctions are competitive, yet only 75.1% (sum of the areas with $EI>0$) are efficient.
@@
We note Fig. 4 shows a somewhat strict definition of competitiveness (i.e., $CI\geq 0$). In reality, builders may bid slightly lower than their true value to ensure profits. Therefore, it is reasonable to relax and consider an auction with $CI\geq-\delta$ to be competitive for a small $\delta$. We call this notion $\delta$-competitive. As shown in Fig. 13, when we consider $\delta=1\%$ or $\delta=2\%$, the percentage of uncompetitive auctions reduces to about 6% and 5%, respectively. That is, under the slightly relaxed definition, approximately 95% auctions were competitive.
In the interest of space, we leave the analysis of efficiency and the implication of MEV on efficiency to Appx. D.1.
@@
MEV’s implication on auction outcomes.   To further understand how the magnitude of MEV in a slot affects the competitiveness and efficiency of the auction, we categorize every slot into one of three tiers according to the winning bid (as a proxy to the significance of the MEV opportunity in that slot): low MEV (below 0.0215 ETH, the 10th percentile of the winning bid values), high MEV (above 0.3933 ETH, 95th percentile) and medium MEV in between. The ratio of competitive auctions in each tier is shown in Fig. 5. We note that competitiveness decreases as the MEV of a slot increases.
@@
Figure 5: The ratio of auctions that are competitive and efficient across various MEV tiers.
Finding I: MEV-Boost auctions are generally competitive, while the competitiveness of MEV-Boost auctions deteriorates as a slot’s MEV increases.
Loss from uncompetiveness.   For an uncompetitive auction, the loss from uncompetitiveness is the difference between the second-highest true value and the actual winning bid value, representing the loss a validator incurs compared to what it would have received in a competitive auction. We computed the loss from uncompetitiveness using our dataset, and the results are shown in Tab. 2. The loss from uncompetitiveness ranges from 0.5% to 1.7% during the periods covered.
Table 2: Losses from uncompetitiveness and inequality.
Time Slots Profits (ETH) Losses-un (ETH) (%) Losses-in (ETH) (%)
April 9-15 28,385 2,704.4 46.9 (1.7) 312.1 (11.5)
May 1-7 30,300 9,331.7 115.8 (1.2) 518.6 (5.6)
June 1-7 35,443 4,341.8 25.1 (0.6) 342.2 (7.9)
July 1-7 36,040 3,938.8 19.1 (0.5) 246.1 (6.3)
August 1-7 17,831 2,135.5 12.5 (0.6) 146.6 (6.9)
September 1-7 35,815 3,939.5 48.5 (1.2) 247.7 (6.3)
October 1-7 39,131 3,427.0 28.9 (0.8) 234.1 (6.8)
November 1-7 39,848 4,736.2 45.6 (1.0) 305.0 (6.4)
December 1-7 38,686 5,884.4 80.0 (1.4) 365.8 (6.2)
4.2 Inequality in Block-building Capacities
The previous section shows that the MEV-Boost auctions are competitive in most cases. However, a proposer may still incur loss due to a significant block-building capacity gap between builders. In MEV Boost auctions, this quantity is captured by the true values of the builder’s bids.
To quantify the disparity of block-building capacity, we compute the quartile coefficient of dispersion (QCD) [88] of the highest true values in a given slot. Specifically, suppose the true values of different builders in slot $s$ are $\vec{tv}_{s}$. Let $Q_{1}(\vec{x})$ and $Q_{3}(\vec{x})$ representing the first and third quartiles of $\vec{x}$ (i.e., 25% and 75% percentiles). The QCD of slot $s$ is
QCD is a real value in $[0,1]$. A larger QCD indicates greater dispersion (the intuition is that in a dataset with big dispersion, $Q_{3}$ is much larger than $Q_{1}$, thus $QCD$ approaches $1$). Since it uses quartiles, QCD is robust to outliers [88].
Evaluation results.   We categorize the builders into three groups based on their market share from April to December 2023: top builders (top 5), middle builders (6-15), and tail builders (16-25). To understand how the magnitude of MEV in a slot affects the block-building capability, we follow the same methodology in Sec. 4.1 and categorize the auctions into three tiers: low MEV, medium MEV, and high MEV.
In Fig. 6, we plot the distribution of the QCD for MEV-Boost auctions from April 2023 to December 2023 for different builder groups and MEV tiers. (We sample the first week of each month. See Sec. 3.1 for details.) It reveals two trends. First, within each MEV tier, there is a clear increase in inequality (a high QCD means high inequality) from the top to the middle and then to the tail. This suggests that while top builders have comparable block-building capacities, middle and tail builders have significantly different capacitiesx. This might be good news: top builders having similar abilities means they can meaningfully compete, which is positive for the decentralization of the builder market. Second, the inequality worsens as the MEV tier increases. One possible reason is that providers may give significant MEV opportunities to select builders, as we will discuss in Sec. 5.3.
Figure 6: Violin graphs for the distribution of the QCD across three MEV tiers, for top, middle, and tail builders. The QCD measures the inequality of block-build capacity across builders. This figure reveals that such inequality worsens as MEV increases. Within a given MEV tier, inequality is greatest among tail builders, followed by middle builders, and is least among top builders.
Finding II: The inequality in block building is lowest among top builders, higher among middle builders, and most pronounced among tail builders, and the inequality worsens as a slot’s MEV increases.
Loss from inequality.   For each auction, the loss from inequality is the difference between the highest and second-highest true values, reflecting the potential loss a validator may face due to unequal block-building capacities, as a competitive auction only ensures the winning bid slightly exceeds the second-highest true value.
As shown in Tab. 2, the loss from inequality ranges from 5.6% to 11.5% during the periods covered by our dataset. Comparing the two types of loss, we find that inequality is the proposers’ primary source of loss. Note that the time range of the full bids is from April to December 2023, and the inequality in the current builder market may worsen (as discussed in Sec. 5.3).
5 What Caused Inequality in Block-building Capacities
Given the significant losses proposers incur due to the inequality in builders’ block-building capacity, we further investigate the causes of this inequality in this section.
Although the builder market is technically open to all, the barrier to accessing profitable private order flows (Sec. 2.2) forms an implicit access barrier.
To illustrate the importance of private order flow to the success of builders, Fig. 7 plots the fraction of builder income from private order flows. Currently, approximately 80% of the block value comes from private order flows.
Figure 7: Fraction of builder income from private order flows.
Finding III: The private order flow is crucial for builders to win the MEV-Boost auction because it contributes to more than 80% of the block value in over 50% of blocks.
Where can new builders get private order flows? The entities who provide (or sell) private order flows to builders are called private order flow providers. Different providers impose different access barriers to their private order flows. Some providers, like MEV Blocker or MEV-Share, are open to new builders who satisfy certain requirements, while some (e.g., most searchers) are almost opaque to the public.
To understand what causes the access barrier for new builders, we first identify important “pivotal” providers and then analyze the requirements to access private order flows from these representative pivotal providers.
5.1 Pivotal Private Order Flow Providers
We now present our method to quantify the importance of individual providers. An existing approach [53, 54] is to compute the contribution of each provider to the builders’ income (Sec. 2.3) during a specific period. However, this approach only considers the magnitude of impact but neglects the duration (in terms of time or the number of auctions). For example, a single private transaction, such as one3330xf0464b01d962f714eee9d4392b2494524d0e10ce3eb3723873afd1346b8b06e4 transferring 457.55 ETH to the builder as a tip, could contribute significantly to the builder’s income, but this provider may only influence the MEV-Boost auction in a single slot. Thus, we assess the significance of a provider as the number of auctions for which they played an important role; in particular, we focus on the providers that can affect half of the MEV-Boost auctions in over two weeks, thereby capturing their sustained impact.
Figure 8: A provider $P$ is pivotal if the winner of an auction could have lost without order flows from $P$. The pivotal level of top-5 pivotal providers over time. A line may not start from the beginning because the provider did not exist at that point. Maestro became an independent provider around Dec 2023.
For each pivotal provider, we compute the fraction of auctions in which it is pivotal. We call this metric the pivotal level, a number from 0 to 1. Fig. 8 plots the pivotal level of the top-5 providers. A builder who lacks access to any of their order flows will lose the majority of auctions. However, we found that order flows from these providers are inaccessible to all builders, which requires further investigation. Other observations from Fig. 8 are provided in Appx. F.2.
5.2 Trust Barrier
A direct reason for the access barrier is that private order flow providers are at a disadvantage when interacting with builders and must trust builders not to unbundle the transactions. However, in practice, builders occasionally unbundle searcher’s bundles. For example, the f1b builder unbundled a bundle from the searcher jaredfromsubway.eth at block height 18476515. Although the searcher was compensated afterward, this incident highlights the fragile security foundation of the MEV-Boost auction.
Due to the lack of trustworthy fair exchange mechanisms, providers have to rely on an informal reputation system for self-protection, sending order flows only to builders who meet their requirements. To understand the access barrier caused by the trust crisis, we investigate different providers and their requirements for accessing order flows.
Searchers.   Since searchers are operated by anonymous entities without any public communication channels, there is no public statement about which builders may access their flows. Existing studies indicate that searchers tend to send their order flow to the top builders [81] or share it exclusively with vertically integrated builders [81, 80, 40].
Telegram trading bots.   The requirements for Telegram trading bots, in terms of sending order flow to builders, depend on whether they utilize channels like MEV Blocker [56, 48]. If so, their requirements are aligned with those of these channels. Banana Gun does not explicitly state the requirements for builders but we observe its preference for Titan, which we will discuss later.
The channels impose various requirements on a builder’s reputation. Market share—the percentage of blocks built by a given builder—is a metric most relay providers use to evaluate builders’ reputations. Intuitively, a builder with a high market share, seen as reputable, is less likely to compromise private order flows as doing so will forgo their future revenue. MeowRPC [58] sends the order flow to trusted builders with high market share. Similarly, MEV-Share [37] requires that a builder possesses a competitive market share. MEV Blocker [60] requires that a builder must maintain at least 1% market share over each week of one month.
Subsidization.   New builders face a “chicken and egg” problem as they need access to private order flows to win auctions and gain market share, but the private order flow providers only serve builders with a decent market share. In practice, builders pay out of pocket to overbid in MEV-Boost auctions, a practice known as block subsidization [36], representing an entry barrier for new builders. We quantify the minimal cost for a new builder to meet the MEV Blocker’s requirements (Appx. F.5), which reflects the costs of building a reputation. The minimal cost continues to increase and can be up to 7.4 ETH. This cost represents a lower bound since we do not have the corresponding numbers for other providers. Note that although subsidization can help meet the requirements set by channels, it cannot ensure that searchers’ trust concerns are addressed, allowing builders to receive their order flows.
5.3 Incentive for Integration
Another access barrier to private order flows is that providers exclusively share order flows with specific builders, known as integration [40]. We compute each provider’s builder-specific pivotal level to understand the relationship between providers and builders. For a given provider $P$, its pivotal level specific to builder $B$ is the pivotal level of the MEV-Boost auctions where $B$ is the winner.
Figure 9: Weekly builder-specific pivotal level for top-3 builders. MEV-Blocker and MEV-Share are hidden for clarity.
Fig. 9 plots selected providers’ builder-specific pivotal levels for the top three builders, ranked based on their weekly market share in July 2024. We include providers whose weekly pivotal levels exceed 25% for any of the three builders over two months to illustrate these builder-specific pivotal levels. MEV Blocker and MEV-Share are excluded from this analysis, as all builders connected to them reportedly receive the same order flows. We note that some providers have a pivotal level exceeding 50% only with specific builders, suggesting a possible collaborative relationship. The searcher 0xA69b… consistently influences over 50% of the MEV-Boost auctions won by Beaver, while the searchers 0x0087…, 0x2800…, and 0x51c7… successively affect those won by Rsync. These pivotal providers are all integrated searchers of the corresponding builders that have been identified in previous works [42, 40]. This observation corroborates findings from previous studies [42, 40, 80], confirming that the order flows from integrated searchers are crucial to those builders.
Furthermore, we observe a clear variation in the importance of providers for different builders. For example, the significance of Banana Gun to Titan is greater than it is to other builders. We further investigate the relationship between Banana Gun and Titan and suppose that Banana Gun only sends valuable order flow to Titan when MEV is significant. A case in this point is slot 8614161, where only Titan received the valuable order flow from Banana Gun. The much lower bids from other builders suggest they didn’t receive this order flow; otherwise, they could have outbid Titan. We illustrate this new integration pattern in Appx. F.3. The results in Fig. 15 indicate that this integration pattern also applies to jaredfromsubway.eth and Beaver. Both integrations emerged around January 2024.
There is a clear incentive for integration between providers and builders. If MEV from a provider can determine the auction outcome, sending order flow to an integrated builder is better than sending it to multiple builders. The integrated builder only needs to bid slightly higher to win, and they can share the surplus MEV. However, this integration is detrimental to proposers, as it exacerbates the inequality in builders’ block-building capacities and can eventually lead to significant losses for them.
Finding IV: Two providers (Banana Gun and jaredfromsubway.eth) exclusively share high-value order flows with specific builders (Titan and Beaver) respectively, suggesting a new integration pattern.
6 Related Work